Find the last two digits of 3^{1000}.   How do I know phi(100) = 2^2 * 5^2 *(1- 1/2)* (1-1/5) ???? This appears and feels it has been done out of thin air. Then, Fermitas ?!?
Answer #1

Question asked:

To prove:

\phi (100)=2^{2}\times 5^{2}\times (1-\frac{1}{2})\times (1-\frac{1}{5})

Proof:

\phi (n) is the number of positive integers \leq n that are relatively prime to n.

\phi (n) is called Euler's totient.

Example :

\phi (9)=6,

   since 1,2,3,4,5,7 and 8 are relatively prime to n.

Formula to find Euler's totient \phi (n):

Step 1:

Express n in terms of its relatively prime numbers as follows:

n=P^{^{a}}_{_{1}}\times P^{^{b}}_{_{2}}\times P^{^{c}}_{_{_{3}}}\times ...

where P1, P2, P3, ... are prime factors of n.

Then,

Euler totient\phi (n) is given by:

\phi (n)=n\times (1-\frac{1}{P_{_{1}}})\times (1-\frac{1}{P_{2}})\times (1-\frac{1}{P_{_{3}}})\times ...   (1)

Here,

n = 100

Expressing n in terms of prime factors, we get:

100 = 22 X 52

Now, applyig formula (1), we get:

\phi (100)=100\times (1-\frac{1}{2})\times (1-\frac{1}{5})

i.e.,

\phi (100)=2^{2}\times 5^{2}\times (1-\frac{1}{2})\times (1-\frac{1}{5})

This is the required result.